OpenStudy (anonymous):

In a table of values for a pattern, P = 12 when n = 3. Determine the equation that might represent the pattern. Select one: a. P = 4(n+6) b. P =4(6 - n) c. P = 4n + 6 d. P = 24 - 3n

OpenStudy (tkhunny):

Have you considered substituting the given value of n into each equation and seeing which produces the given value for P?

OpenStudy (anonymous):

hmm no not yet

OpenStudy (anonymous):

idk..

OpenStudy (tkhunny):

Well, do it! a. P = 4(n+6) 4(3+6) = 4*9 = 18 and that is not 12 Try the next one.

OpenStudy (anonymous):

ohhhh ok